# Scrambling In Charged Hairy Black Holes and the Kasner Interior

### v2

@@ II Interior of Charged Hairy Black Hole
$\Phi^{\prime\prime}+\Phi^{\prime}\frac{\quantity((\gamma\phi^{2}+1)\chi^{\prime}+4\phi\phi^{\prime})}{(1+\gamma\phi^{2})}-\Phi\frac{2q^{2}\phi^{2}}{r^{2}f(1+\gamma\phi^{2})}$ $=0,$ (11)
@@
These equations are coupled and highly non-linear. Therefore, we seek numerical calculations for solving the equations of motion. We also set the AdS radius to $L=1$ and the horizon radius to $r_{h}=1$ through the calculations.
With vanishing scalar field $\phi$, the solution is given by the axion-Reissner-Nordström (aRN) black hole, with
$f(r)$ $=1-\frac{r^{3}}{r_{h}^{3}}-\frac{r^{2}\zeta^{2}}{2}+\frac{r^{3}\zeta^{2}}{2r_{h}}+\frac{r^{4}\rho^{2}}{4}-\frac{r^{3}r_{h}\rho^{2}}{4},$ (14)
[…]
$\Phi(r)$ $=\mu-\rho r.$ (16)
@@ II.2 Flow Through General Kasner Interior
×$\times(\partial_{\nu}\phi^{*}+iqA_{\nu}\phi^{*})-\frac{1}{2}m^{2}|\phi|^{2}\bigg{)}$
@@ III Mutual Information and Lyapunov Exponent
Figure 3: Penrose diagram of the black hole spacetime perturbed by gravitational shock waves $\alpha$. $r_{c}$ represents the critical turning point while the horizontal red line represents the minimal surface. $I,II,$ and $III$ are the segments that we are using in calculating the minimal surface.
@@ III.1 Holographic Calculation of Mutual Information
Here, $L_{y}$ represents the total length of the $y$-direction, while $r_{t}$ is the turning point in the bulk where $\dot{r}=0$. The dot $\dot{}$ indicates differentiation with respect to $t$, whereas the prime denotes differentiation with respect to $r$. The entanglement entropy $S_{A\cup B}$ is calculated by the sum of the surface areas located at $x=0$ and $x=l_{x}$.
@@
is the tortoise coordinate, following [12, 18, 19]. The first segment is an integration from the boundary with $(U,V)=(1,-1)$ to the horizon with $(U,V)=(U_{1},0)$. The second segment stretches from the horizon with $(U,V)=(U_{1},0)$ to the turning point at $r=r_{t}$ denoted by $(U,V)=(U_{2},V_{2})$. The third segment now runs from the turning point at $(U,V)=(U_{2},V_{2})$ to the location deformed by the shock waves $\alpha$ at the horizon with $(U,V)=(0,\alpha/2)$. These three segments are depicted in Figure 3.
@@ III.2 Lyapunov Exponent from Boundary Deformations
$-\frac{4}{r_{c}^{5}}+\frac{\zeta^{2}}{r_{c}^{3}}+\frac{\rho^{2}r_{h}^{4}-2\zeta^{2}r_{h}^{2}+4}{4r_{c}^{2}r_{h}^{3}}=0.$ (63)
@@ III.3 Near-Singularity Area Functional
[…]
@@ V Summary and Discussions
An interesting future direction is to study how rotating shock waves [16], or even rotating and charged shock waves [19], disrupt the interior structure of rotating and charged hairy black holes [39]. Understanding how these perturbations affect the black hole interior could provide new insights into the role of rotation and charge in chaotic dynamics. Furthermore, we are also interested in investigating how this relevant deformation influences the traversability of a wormhole induced by double-trace deformations on the boundary [40, 41]. This is particularly intriguing because information passing through a traversable wormhole necessarily probes the black hole interior before emerging in another region. Additionally, recent work on analytical solutions of scalar-hairy black holes [42] provides a foundation for further studies of chaotic behavior in specific limits, such as the low-temperature limit $T\to 0$. We hope this finding can help us in investigating the chaotic behavior especially in some certain limit of hairy black holes. We leave these investigations in future works.

### v3

@@ II Interior of Charged Hairy Black Hole
$\Phi^{\prime\prime}+\Phi^{\prime}\frac{\quantity((\gamma\phi^{2}+1)\chi^{\prime}+4\phi\phi^{\prime})}{2(1+\gamma\phi^{2})}-\Phi\frac{2q^{2}\phi^{2}}{r^{2}f(1+\gamma\phi^{2})}$ $=0,$ (11)
@@ II.2 Flow Through General Kasner Interior
$\times(\partial_{\nu}\phi^{*}+iqA_{\nu}\phi^{*})-\frac{1}{2}m^{2}|\phi|^{2}\bigg{)}$
@@ III Mutual Information and Lyapunov Exponent
Figure 3: Penrose diagram of the black hole spacetime perturbed by gravitational shock waves $\alpha$. $r_{c}$ represents the critical turning point while the horizontal red line represents the minimal surface. $I,I$,II, and $III$ are the segments that we are using in calculating the minimal surface.
@@ III.1 Holographic Calculation of Mutual Information
[…]
@@
is the tortoise coordinate, following [12, 18, 19]. The first segment is an integration from the boundary with (U,V)=(1,1)(U,V)=(1,-1) to the horizon with $(U,V)=(U_{1},0)$. The second segment stretches from the horizon with $(U,V)=(U_{1},0)$ to the turning point at $r=r_{t}$ denoted by $(U,V)=(U_{2},V_{2})$. The third segment now runs from the turning point at $(U,V)=(U_{2},V_{2})$ to the location deformed by the shock waves $\alpha$ at the horizon with $(U,V)=(0,\alpha/2)$. These three segments are depicted in Figure 3.
@@ III.2 Lyapunov Exponent from Boundary Deformations
$\frac{4}{r_{c}^{5}}+\frac{\zeta^{2}}{r_{c}^{3}}+\frac{\rho^{2}r_{h}^{4}-2\zeta^{2}r_{h}^{2}+4}{4r_{c}^{2}r_{h}^{3}}=0.$ (60)
@@ III.3 Near-Singularity Area Functional
$\mathcal{A}=L_{y}\int dr\frac{1}{\sqrt{-f_{K1}}r^{(7+2c^{2})/2}}\quantity[1-\quantity(\frac{K^{2}}{f_{K1}e^{-\chi_{K1}}})r^{1+c^{2}}]^{1/2},$ (63)
@@ V Summary and Discussions
It has been shown [23, 21, 39] that the charged hairy black hole does not have an inner horizon. On the other hand, it is also known that the charged black hole solution (with $\phi=0$) has inner horizon instability due to mass inflation (see, for example, [40, 41, 42] and references therein). It is worth noting that the absence of an inner horizon in the charged hairy black hole geometry might be related to the suppression of the mass inflation effect. Although our numerical results for chaos-related quantities such as the Lyapunov exponent smoothly approach those of the axion-Reissner-Nordström solution in the small-$\phi_{0}$ limit, the connection between this limit and the suppression of mass inflation near the inner horizon remains unclear. We believe that future work in this direction could shed important light on the problem. Several earlier studies (for example, [21,39]) have analytically shown that solutions with $f(r_{I})=0$, where $r_{I}$ denotes the would-be inner horizon, do not arise in charged hairy black holes. These works could serve as a starting point for investigating the relation between the absence of the inner horizon and the suppression of mass inflation.
In this work, we restrict our analysis to $L=1$ case for numerical simplicity, where $L$ is the AdS radius. However, learning about the large-$L$ limit is also interesting since it might coincide with the asymptotically flat black holes. In our previous work [19], we see that when the AdS radius becomes large, the minimal instantaneous Lyapunov exponent approaches its surface gravity, thus saturating the Maldacena-Shenker-Stanford bound. In our current setup, the numerical calculations fail when the AdS radius is dramatically increased. However, we may find interesting results if we could study the solution of the hairy black hole under the large-$L$ expansion and compare the result to chaos in flat-space holography [43]. Asymptotically flat black holes are of particular interest due to their relevance to astrophysical observations. Although still speculative, certain chaotic properties of black holes such as the Lyapunov exponent derived from particle orbits may have connections to observable features like black hole imaging [44]. Therefore, investigating the relation between the flat-space limit of holographic calculations and the chaotic behavior of particle motion around black holes could provide valuable insights for future studies.
Another interesting future direction is to study how rotating shock waves [16], or even rotating and charged shock waves [19], disrupt the interior structure of rotating and charged hairy black holes [45]. Understanding how these perturbations affect the black hole interior could provide new insights into the role of rotation and charge in chaotic dynamics. Furthermore, we are also interested in investigating how this relevant deformation influences the traversability of a wormhole induced by double-trace deformations on the boundary [46, 47]. Additionally, recent work on analytical solutions of scalar-hairy black holes [48] provides a foundation for further studies of chaotic behavior in specific limits, such as the low-temperature limit $T\to 0$. We hope this finding can help us in investigating the chaotic behavior especially in some certain limit of hairy black holes. We leave these investigations in future works.